# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 5 Deift–Zhou steepest descent analysis
  1. 1.
    $s\in\mathcal{K}_{out}$, where ${\mathcal{K}_{out}}$ is a closed subset contained outside of $EoT$.
  2. 2.
    $s\in\mathcal{K}_{in,+}$, where ${\mathcal{K}_{in,+}}$ is a compact subset of $\{\mathrm{Re}\,(s)>0\}\cap EoT$.
  3. 3.
    $s\in\mathcal{K}_{in,-}$, where ${\mathcal{K}_{in,-}}$ is a compact subset of $\{\mathrm{Re}\,(s)<0\}\cap EoT$.
Note that for $s\in i{\mathbb{R}}\cap EoT$ the elliptic curve is degenerate, namely, two branchpoints coincide and the curve becomes of genus zero. Here a different analysis would be needed, analogous to the one needed when $s$ belongs to the boundary of $EoT$. Near the corners $s=\pm\frac{i}{2}$ a yet different analysis would be needed, which involves the construction of special local parametrices. We do not discuss these transitional regions.
@@
Theorem 5.1
For any closed subset $\mathcal{K}$ on the outside of $EoT$, the RHP 2.1 (with $t=ns$) is solvable for $n$ sufficiently large, and hence the poles of the rational solutions must be inside a neighbourhood of $EoT$.
The Deift-Zhou steepest descent analysis, formalized in [27] and in many articles thereafter requires a number of transformations of the problem into equivalent ones. We describe briefly below these problems. We recall that our starting point is the following
Problem 5.1
@@
$Y(z_{+})=Y(z_{-})\begin{bmatrix}1&z^{{{K}}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix},\qquad z\in\gamma$ (5.1)
$Y(z)=\left(\mathbf{1}+\mathcal{O}(z^{-1})\right)z^{n\sigma_{3}},\ \ \ z\to\infty.$ (5.2)
@@
where the domain is ${\mathbb{C}}\setminus\Gamma_{0}^{1}$, where $\Gamma_{0}^{1}$ denotes an arc homotopic to $[0,1]$ at fixed endpoints. For $s$ outside $EoT$ we shall choose $\Gamma_{0}^{1}=[0,1]$ (the segment).
It will be necessary for the analysis inside $EoT$ to partially homotopically retract the integration contour $\gamma$ along a subarc of $\Gamma_{0}^{1}$ (from the left and right of it): in that case the RHP (5.1) will take a slightly different form due to the fact that the function $Q$ (5.3) has a jump discontinuity such that $Q(z_{+})=Q(z_{-}){\rm e}^{2i\pi\rho}$. In this case the jump matrix of the RHP (5.1) along such a subarc of $\Gamma_{0}^{1}$ needs to be replaced by
$\begin{bmatrix}1&Q(z_{+})\left({\rm e}^{-2i\pi\rho}-1\right){\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix}.$ (5.4)
5.1 Asymptotic analysis for $s$ outside $EoT$
@@
A direct verification shows that $W$ solves the following
Problem 5.2
@@
$W(z_{+})=W(z_{-})J_{W}(z),\ \ \ \ z\in\gamma$ (5.6)
$W(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\ \ \ \ z\to\infty,$ (5.7)
@@
$J_{W}(z)=\begin{bmatrix}{\rm e}^{-n(g(z_{+})-g(z_{-}))}&Q(z){\rm e}^{n\left(\frac{s}{z}+g(z_{+})+g(z_{-})+\ell\right)}\\ 0&{\rm e}^{n(g(z_{+})-g(z_{-}))}\end{bmatrix}=\begin{bmatrix}{\rm e}^{-\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}&Q(z){\rm e}^{\frac{n}{2}\left(\varphi(z_{+})+\varphi(z_{-})\right)}\\ 0&{\rm e}^{\frac{n}{2}\left(\varphi(z_{+})-\varphi(z_{-})\right)}\end{bmatrix}.$ (5.8)
Lens Opening: $W\to T$.
Refer to Fig. 5. The lens regions $\Lambda_{\pm}$ are the two regions bounded by $\Gamma_{m}$ and the arcs $\mathscr{L}_{\pm}$ chosen arbitrarily in the regions where $\mathrm{Re}\,\varphi>0$. The process of “opening the lenses” consists in re-defining the matrix $W$ within those regions. We thus define
$T_{0}(z):=\left\{\begin{array}[]{lc}W(z)&z\not\in\Lambda_{\pm}\\ \displaystyle W(z)\begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\Lambda_{-}\\[10.0pt] W(z)\begin{bmatrix}1&0\\ -\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\Lambda_{+}.\end{array}\right.$ (5.9)
A direct computation shows that the matrix $T_{0}(z)$ satisfies the following RHP
Problem 5.3
@@
$T_{0}(z_{+})=T_{0}(z_{-})J_{T_{0}}(z),\ \ \ \ z\in\gamma$ (5.10)
$T_{0}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\ \ \ \ z\to\infty$ (5.11)
@@
$J_{T_{0}}(z)=\left\{\begin{array}[]{lc}J_{{}_{W}}(z)&z\in\Gamma_{c}\\[12.0pt] \begin{bmatrix}1&0\\ \frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix}&z\in\mathscr{L}_{\pm}\\[15.0pt] \begin{bmatrix}0&Q(z)\\ -\frac{1}{Q(z)}&0\end{bmatrix}&z\in\Gamma_{m}.\end{array}\right.$ (5.12)
At this point we have obtained a RHP where the jump matrices on $\Gamma_{c}\cup\mathscr{L}_{\pm}$ converge pointwise in the relative interior to the identity matrix, but not uniformly.
To further normalize the problem we need to construct the outer parametrix, namely the (explicit) solution of an auxiliary RHP where we simply drop the jump conditions on $\Gamma_{c}\cup\mathscr{L}_{\pm}$.
Outer parametrix.
We seek the solution of the following “model problem”
Problem 5.4
@@
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z)\\ -\frac{1}{Q(z)}&0\end{bmatrix},\qquad z\in\Gamma_{m}$ (5.13)
$M_{{}_{Q}}(z)=\mathbf{1}+\mathcal{O}(z^{-1}),\qquad|z|\to\infty$ (5.14)
$M_{{}_{Q}}(z)=\mathcal{O}\left(\frac{1}{\left(z\mp\frac{is}{2}\right)^{\frac{1}{4}}}\right),\qquad z\to\pm\frac{is}{2}.$ (5.15)
We now solve RHP 5.4. To this end we need to construct a special scalar function $S(z)$, often called the Szegö function, satisfying the following scalar boundary value problem
$S(z_{+})+S(z_{-})=\ln Q(z),\ \ \ \ z\in\Gamma_{m}$ (5.16)
$\sup_{z\in{\mathbb{C}}\setminus\Gamma_{m}}|S(z)|<+\infty.$ (5.17)
The solution is given by the following expression
$S(z)=R(z)\int_{\Gamma_{m}}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2i\pi},\qquad R(z):=\sqrt{z^{2}+\frac{s^{2}}{4}},$ (5.18)
@@
We can actually simplify the Szegö function by a contour deformation recalling that $\ln Q(w_{+})=\ln Q(w_{-})+2i\pi{{\rho}}$ for $z$ on the segment $[0,1]$; indeed by a contour deformation we can rewrite $S(z)$ as follows
$S(z)=$ $R(z)\int_{\Gamma_{m}}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2i\pi}$ (5.19)
$=$ $-\frac{R(z)}{2}\oint_{\odot\Gamma_{m}}\frac{\ln Q(w){\mathrm{d}}w}{R(w)(w-z)2i\pi},$ (5.20)
where the symbol $\odot\Gamma_{m}$ stands for a counterclockwise loop leaving $\Gamma_{m}$ in its interior region and the segment $[0,1]$ on the exterior. Using then Cauchy’s residue theorem we get
$S(z)=$ $\frac{1}{2}\ln Q(z)-\frac{R(z)}{2}\int_{0}^{1}\frac{\left(\ln Q(w_{+})-\ln Q(w_{-})\right){\mathrm{d}}w}{(w-z)R(w)2i\pi}=$ (5.21)
$=$ $\frac{\ln Q(z)}{2}-\frac{R(z)}{2}\int_{0}^{1}\frac{{{\rho}}{\mathrm{d}}w}{R(w)(w-z)}-\frac{R(z)}{2}\int_{-\infty}^{0}\frac{{{K}}{\mathrm{d}}w}{R(w)(w-z)}=$
$=$ $\frac{{K}}{2}\ln\left(\frac{\left(R(z)-z\right)\left(R(0)+R(z)\right)}{s/2}\right)+\frac{{\rho}}{2}\ln\left({\frac{z+\frac{s^{2}}{4}+R(z)R(1)}{\frac{s}{2}\,\left(R(z)+R(0)\right)}}\right).$ (5.22)
The simplest way to verify this latest formula is to verify $S(z_{+})+S(z_{-})=\ln Q(z)$ using that $R(z_{+})=-R(z_{-})$; also one needs to verify that the expression has no singularities except the discontinuity across the branch cut of $R(z)$, and that it is bounded at infinity. We leave the instructive but tedious verification to the reader.
Remark 5.1
@@
$S(\infty)=\frac{{K}}{2}\ln\left(\frac{s}{4}\right)+\frac{{\rho}}{2}\ln\left(\frac{2}{s}+\sqrt{1+\frac{4}{s^{2}}}\right).$ (5.23)
With these preparations we state
Proposition 5.1 (Solution of RHP 5.4)
The solution of RHP 5.4 is given by
$M_{{}_{Q}}(z)$ $:={\rm e}^{-S(\infty)\sigma_{3}}M(z){\rm e}^{S(z)\sigma_{3}}$ (5.24)
$M(z)$ $:=\left(\frac{z-\frac{is}{2}}{z+\frac{is}{2}}\right)^{\frac{\sigma_{2}}{4}}=\frac{1}{2}\begin{bmatrix}{F}+\frac{1}{F}&-i\left({F}-\frac{1}{F}\right)\\ i\left({F}-\frac{1}{F}\right)&{F}+\frac{1}{F}\end{bmatrix}$ (5.25)
${F}$ $:=\left(\frac{z-\frac{is}{2}}{z+\frac{is}{2}}\right)^{\frac{1}{4}}$ (5.26)
where the branch cut of ${F}$ runs along $\Gamma_{m}$ and the determination is chosen so that ${F}(z)\to 1$ as $|z|\to\infty$.
@@ 5.1.1 Conclusion of the steepest descent analysis
The final steps of the analysis requires the construction of local parametrices near the endpoints ±$\pm\frac{is}{2}$. Namely, one fixes two disks $\mathbb{D}_{\pm}$ centered at each of the two points; this part is a quite standard construction and therefore we only sketch the main points. Let $\mathbb{D}_{\pm}$ be two disjoint disks centered at ±$\pm\frac{is}{2}$ (respectively) and of radii, say, $r=\frac{|s|}{4}$ (the radius is not important for the discussion as long as they are small enough so as not to contain the origin):
$\mathbb{D}_{\pm}:=\left\{\left|z\mp\frac{is}{2}\right|<\frac{|s|}{4}\right\}.$ (5.27)
In a neighbourhood of $a_{\pm}=\frac{\pm is}{2}$ the effective potential $\varphi$ (4.18) has the following behaviour
$\varphi(z;s)=\frac{2^{\frac{5}{2}}}{3}\left(\frac{z}{a_{\pm}}-1\right)^{\frac{3}{2}}\left(1+\mathcal{O}(z-a_{\pm})\right)\mod 2i\pi,\qquad z\sim a_{\pm}.$ (5.28)
In fact the additive constant is 00 in the case of $a_{-}$ and $-2i\pi$ for $a_{+}$, but since $\varphi$ appears always in the exponent, this is irrelevant.
@@
$-\frac{4}{3}\zeta_{-}^{\frac{3}{2}}$ $=n\varphi(z)=n\frac{2^{\frac{5}{2}}}{3}\left(\frac{z}{a_{-}}-1\right)^{\frac{3}{2}}\left(1+\mathcal{O}(z-a_{-})\right)$ (5.29)
$-\frac{4}{3}\zeta_{+}^{\frac{3}{2}}$ $=n\left(\varphi(z)+2i\pi\right)=n\frac{2^{\frac{5}{2}}}{3}\left(\frac{z}{a_{+}}-1\right)^{\frac{3}{2}}\left(1+\mathcal{O}(z-a_{+})\right).$ (5.30)
From the above formulas it appears that both $\zeta_{\pm}$ define a conformal map from the two disks $\mathbb{D}_{\pm}$ (respectively) to a neighbourhood $\mathscr{D}_{n}$ of the origin which is homothetically expanding with $n^{\frac{3}{2}}$. The determination of the fractional root can be chosen so that $\zeta_{\pm}$ map the main arc $\Gamma_{m}$ (where $\mathrm{Re}\,\varphi=0$ and $\mathrm{Im}\,\varphi(z_{+})$ is decreasing from 00 at $a_{-}$ to $-2i\pi$ at $a_{+}$) to the negative real axis in the respective $\zeta_{\pm}$ planes, with the points $a_{\pm}$ being mapped to $\zeta_{\pm}=0$. The arc $\Gamma_{c}$ can be chosen so that it is mapped to the positive $\zeta_{\pm}$-axis while the arcs of ±$\mathscr{L}_{\pm}\cap\mathbb{D}_{\pm}$ are mapped to two straight segments with slopes ±$\pm 3\pi/2$.
Then, the jump matrices $J_{T_{0}}$ in the RHP 5.3 restricted to $\mathbb{D}_{+}$ can be rewritten as (we focus on the case of $\mathbb{D}_{+}$ for definiteness)
$J_{T_{0}}(z)=\left\{\begin{array}[]{lc}\begin{bmatrix}1&Q(z){\rm e}^{-\frac{4}{3}\zeta_{+}^{\frac{3}{2}}}\\ 0&1\end{bmatrix}&z\in\Gamma_{c}\cap\mathbb{D}_{+}\\[12.0pt] \begin{bmatrix}1&0\\ \frac{\displaystyle{\rm e}^{\frac{4}{3}{\zeta_{+}}^{\frac{3}{2}}}}{Q(z)}&1\end{bmatrix}&z\in\mathscr{L}_{\pm}\cap\mathbb{D}_{+}\\[15.0pt] \begin{bmatrix}0&Q(z)\\ -\frac{1}{Q(z)}&0\end{bmatrix}&z\in\Gamma_{m}\cap\mathbb{D}_{+}\end{array}\right..$ (5.31)
@@
$\mathcal{P}_{+}(z):={\bf A}\big{(}\zeta_{+}(z)\big{)}\,Q(z)^{-\frac{\sigma_{3}}{2}}{\rm e}^{\frac{2}{3}\zeta_{+}(z)^{\frac{3}{2}}\sigma_{3}}$ (5.32)
exhibits discontinuities across $(\Gamma_{c}\cup\mathscr{L}_{+}\cup\mathscr{L}_{-}\mathscr{\Gamma}_{m})\cap\mathbb{D}_{+}$ with jump matrices given exactly by (5.31), as a consequence of the RHP satisfied by the matrix ${\bf A}(\zeta)$ shown on Fig. 11. Completely analogous expressions hold for $\mathcal{P}_{-}(z)$.

### v3

@@ 5 Deift–Zhou steepest descent analysis
  1. (1)
    $s\in\mathcal{K}_{\rm out}$, where $\mathcal{K}_{\rm out}$} is a closed subset contained outside of $\rm EoT$},
  2. (2)
    $s\in\mathcal{K}_{{\rm in},+}$, where $\mathcal{K}_{{\rm in},+}$} is a compact subset of $\{\operatorname{Re}(s)>0\}\cap{\rm EoT}$,
  3. (3)
    $s\in\mathcal{K}_{{\rm in},-}$, where $\mathcal{K}_{{\rm in},-}$} is a compact subset of $\{\operatorname{Re}(s)<0\}\cap{\rm EoT}$.
Note that for $s\in{\rm i}{\mathbb{R}}\cap{\rm EoT}$ the elliptic curve is degenerate, namely, two branch-points coincide and the curve becomes of genus zero. Here a different analysis would be needed, analogous to the one needed when $s$ belongs to the boundary of $\rm EoT$}. Near the corners $s=\pm\frac{\rm i}{2}$ a yet different analysis would be needed, which involves the construction of special local parametrices. We do not discuss these transitional regions.
@@
Theorem 5.1.
For any closed subset $\mathcal{K}$ on the outside of $\rm EoT$}, the Riemann–Hilbert Problem 2.4 ((with $t=ns)$ is solvable for $n$ sufficiently large, and hence the poles of the rational solutions must be inside a neighbourhood of $\rm EoT$}.
The Deift–Zhou steepest descent analysis, formalized in [27] and in many articles thereafter requires a number of transformations of the problem into equivalent ones. We describe briefly below these problems. We recall that our starting point is the following.
Problem 5.2.
@@
$Y(z_{+})=Y(z_{-})\begin{bmatrix}1&z^{{{K}}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix},\qquad z\in\gamma,$
$Y(z)=\left(\mathbf{1}+\mathcal{O}\bigl(z^{-1}\bigr)\right)z^{n\sigma_{3}},\qquad z\to\infty.$
@@
where the domain is ${\mathbb{C}}\setminus\Gamma_{0}^{1}$, where $\Gamma_{0}^{1}$ denotes an arc homotopic to $[0,1]$ at fixed endpoints. For $s$ outside $\rm EoT$}, we shall choose $\Gamma_{0}^{1}=[0,1]$ (the segment).
It will be necessary for the analysis inside $\rm EoT$} to partially homotopically retract the integration contour $\gamma$ along a subarc of $\Gamma_{0}^{1}$ (from the left and right of it). In that case, the Riemann–Hilbert Problem 5.2 will take a slightly different form due to the fact that the function $Q$ (5.1) has a jump discontinuity such that $Q(z_{+})=Q(z_{-}){\rm e}^{2{\rm i}\pi\rho}$. In this case, the jump matrix of the Riemann–Hilbert Problem 5.2 along such a subarc of $\Gamma_{0}^{1}$ needs to be replaced by
$\begin{bmatrix}1&Q(z_{+})\left({\rm e}^{-2{\rm i}\pi\rho}-1\right){\rm e}^{\frac{ns}{z}}\\ 0&1\end{bmatrix}$.
5.1 Asymptotic analysis for $\boldsymbol{s}$ outside $\boldsymbol{{\rm EoT}}$
@@
A direct verification shows that $W$ solves the following.
Problem 5.3.
@@
$W(z_{+})=W(z_{-})J_{W}(z),\qquad z\in\gamma,\qquad W(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\qquad z\to\infty,$
@@
$J_{W}(z)$ $=\begin{bmatrix}{\rm e}^{-n(g(z_{+})-g(z_{-}))}&Q(z){\rm e}^{n(\frac{s}{z}+g(z_{+})+g(z_{-})+\ell)}\\ 0&{\rm e}^{n(g(z_{+})-g(z_{-}))}\end{bmatrix}$
$=\begin{bmatrix}{\rm e}^{-\frac{n}{2}(\varphi(z_{+})-\varphi(z_{-}))}&Q(z){\rm e}^{\frac{n}{2}(\varphi(z_{+})+\varphi(z_{-}))}\\ 0&{\rm e}^{\frac{n}{2}(\varphi(z_{+})-\varphi(z_{-}))}\end{bmatrix}$.
Lens Opening: $\boldsymbol{W\to T}$. Refer to Figure 5. The lens regions $\Lambda_{\pm}$ are the two regions bounded by $\Gamma_{m}$ and the arcs $\mathscr{L}_{\pm}$ chosen arbitrarily in the regions where $\operatorname{Re}\varphi>0$. The process of “opening the lenses” consists in re-defining the matrix $W$ within those regions. We thus define
$T_{0}(z):=\begin{cases}W(z),&z\not\in\Lambda_{\pm},\\ \displaystyle W(z)\begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\Lambda_{-},\vskip 2.84526pt\\ W(z)\begin{bmatrix}1&0\\ \displaystyle-\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\Lambda_{+}.\end{cases}$ (5.3)
A direct computation shows that the matrix $T_{0}(z)$ satisfies the following Riemann–Hilbert problem.
Problem 5.4.
@@
$T_{0}(z_{+})=T_{0}(z_{-})J_{T_{0}}(z),\qquad z\in\gamma,\qquad T_{0}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{z}\right),\qquad z\to\infty,$
@@
$J_{T_{0}}(z)=\begin{cases}J_{{}_{W}}(z),&z\in\Gamma_{c},\\ \begin{bmatrix}1&0\\ \displaystyle\frac{{\rm e}^{-n\varphi(z)}}{Q(z)}&1\end{bmatrix},&z\in\mathscr{L}_{\pm},\vskip 2.84526pt\\ \begin{bmatrix}0&Q(z)\\ \displaystyle-\frac{1}{Q(z)}&0\end{bmatrix},&z\in\Gamma_{m}.\end{cases}$
At this point, we have obtained a RHP, where the jump matrices on $\Gamma_{c}\cup\mathscr{L}_{\pm}$ converge pointwise in the relative interior to the identity matrix, but not uniformly.
To further normalize the problem, we need to construct the outer parametrix, namely the (explicit) solution of an auxiliary RHP where we simply drop the jump conditions on $\Gamma_{c}\cup\mathscr{L}_{\pm}$.
Outer parametrix. We seek the solution of the following “model problem”.
Problem 5.5.
@@
$M_{{}_{Q}}(z_{+})=M_{{}_{Q}}(z_{-})\begin{bmatrix}0&Q(z)\\ -\frac{1}{Q(z)}&0\end{bmatrix},\qquad z\in\Gamma_{m},\qquad M_{{}_{Q}}(z)=\mathbf{1}+\mathcal{O}\bigl(z^{-1}\bigr),\qquad|z|\to\infty,$
$M_{{}_{Q}}(z)=\mathcal{O}\left(\frac{1}{\left(z\mp\frac{{\rm i}s}{2}\right)^{\frac{1}{4}}}\right),\qquad z\to\pm\frac{{\rm i}s}{2}.$
We now solve Riemann–Hilbert Problem 5.5. To this end, we need to construct a special scalar function $S(z)$, often called the Szegő function, satisfying the following scalar boundary value problem $S(z_{+})+S(z_{-})=\ln Q(z)$, $z\in\Gamma_{m}$, $\sup_{z\in{\mathbb{C}}\setminus\Gamma_{m}}|S(z)|<+\infty$. The solution is given by the following expression:
$S(z)=R(z)\int_{\Gamma_{m}}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi},\qquad R(z):=\sqrt{z^{2}+\frac{s^{2}}{4}},$ (5.4)
@@
We can actually simplify the Szegő function by a contour deformation recalling that $\ln Q(w_{+})\allowbreak=\ln Q(w_{-})+2{\rm i}\pi{{\rho}}$ for $z$ on the segment $[0,1]$. Indeed, by a contour deformation, we can rewrite $S(z)$ as follows:
$S(z)=R(z)\int_{\Gamma_{m}}\frac{\ln Q(w){\mathrm{d}}w}{R(w_{+})(w-z)2{\rm i}\pi}=-\frac{R(z)}{2}\oint_{\odot\Gamma_{m}}\frac{\ln Q(w){\mathrm{d}}w}{R(w)(w-z)2{\rm i}\pi},$
where the symbol $\odot\Gamma_{m}$ stands for a counterclockwise loop leaving $\Gamma_{m}$ in its interior region and the segment $[0,1]$ on the exterior. Using then Cauchy’s residue theorem, we get
$S(z)={}$ $\frac{1}{2}\ln Q(z)-\frac{R(z)}{2}\int_{0}^{1}\frac{\left(\ln Q(w_{+})-\ln Q(w_{-})\right){\mathrm{d}}w}{(w-z)R(w)2{\rm i}\pi}$
$={}$ $\frac{\ln Q(z)}{2}-\frac{R(z)}{2}\int_{0}^{1}\frac{{{\rho}}{\mathrm{d}}w}{R(w)(w-z)}-\frac{R(z)}{2}\int_{-\infty}^{0}\frac{{{K}}{\mathrm{d}}w}{R(w)(w-z)}$
$={}$ $\frac{{K}}{2}\ln\left(\frac{\left(R(z)-z\right)\left(R(0)+R(z)\right)}{s/2}\right)+\frac{{\rho}}{2}\ln\left({\frac{z+\frac{s^{2}}{4}+R(z)R(1)}{\frac{s}{2}(R(z)+R(0))}}\right).$ (5.5)
The simplest way to verify this latest formula is to verify $S(z_{+})+S(z_{-})=\ln Q(z)$ using that $R(z_{+})=-R(z_{-})$. Also one needs to verify that the expression has no singularities except the discontinuity across the branch cut of $R(z)$, and that it is bounded at infinity. We leave the instructive but tedious verification to the reader.
Remark 5.6.
@@
$S(\infty)=\frac{{K}}{2}\ln\left(\frac{s}{4}\right)+\frac{{\rho}}{2}\ln\left(\frac{2}{s}+\sqrt{1+\frac{4}{s^{2}}}\right).$
With these preparations, we state the following.
Proposition 5.7 (solution of Riemann–Hilbert Problem 5.5).
The solution of Riemann–Hilbert Problem 5.5 is given by
$M_{{}_{Q}}(z):={\rm e}^{-S(\infty)\sigma_{3}}M(z){\rm e}^{S(z)\sigma_{3}},$ (5.6)
$M(z):=\left(\frac{z-\frac{{\rm i}s}{2}}{z+\frac{{\rm i}s}{2}}\right)^{\frac{\sigma_{2}}{4}}=\frac{1}{2}\begin{bmatrix}{F}+\frac{1}{F}&-{\rm i}\left({F}-\frac{1}{F}\right)\\ {\rm i}\left({F}-\frac{1}{F}\right)&{F}+\frac{1}{F}\end{bmatrix},\qquad{F}:=\left(\frac{z-\frac{{\rm i}s}{2}}{z+\frac{{\rm i}s}{2}}\right)^{\frac{1}{4}},$
where the branch cut of F{F} runs along $\Gamma_{m}$ and the determination is chosen so that ${F}(z)\to 1$ as $|z|\to\infty$}.
@@ 5.1.1 Conclusion of the steepest descent analysis
The final steps of the analysis requires the construction of local parametrices near the endpoints $\pm\frac{{\rm i}s}{2}$. Namely, one fixes two disks $\mathbb{D}_{\pm}$ centered at each of the two points. This part is a quite standard construction and therefore we only sketch the main points. Let $\mathbb{D}_{\pm}$ be two disjoint disks centered at $\pm\frac{{\rm i}s}{2}$ (respectively) and of radii, say, $r=\frac{|s|}{4}$ (the radius is not important for the discussion as long as they are small enough so as not to contain the origin),
$\mathbb{D}_{\pm}:=\left\{\left|z\mp\frac{{\rm i}s}{2}\right|<\frac{|s|}{4}\right\}.$
In a neighbourhood of $a_{\pm}=\frac{\pm{\rm i}s}{2}$, the effective potential $\varphi$ (4.4) has the following behaviour:
$\varphi(z;s)=\frac{2^{\frac{5}{2}}}{3}\left(\frac{z}{a_{\pm}}-1\right)^{\frac{3}{2}}\left(1+\mathcal{O}(z-a_{\pm})\right)\mod 2{\rm i}\pi,\qquad z\sim a_{\pm}.$
In fact, the additive constant is 0 in the case of $a_{-}$ and $2{\rm i}\pi$ for $a_{+}$, but since $\varphi$ appears always in the exponent, this is irrelevant.
@@
$-\frac{4}{3}\zeta_{-}^{\frac{3}{2}}=n\varphi(z)=n\frac{2^{\frac{5}{2}}}{3}\left(\frac{z}{a_{-}}-1\right)^{\frac{3}{2}}\left(1+\mathcal{O}(z-a_{-})\right),$
$-\frac{4}{3}\zeta_{+}^{\frac{3}{2}}=n\left(\varphi(z)+2{\rm i}\pi\right)=n\frac{2^{\frac{5}{2}}}{3}\left(\frac{z}{a_{+}}-1\right)^{\frac{3}{2}}\left(1+\mathcal{O}(z-a_{+})\right).$
From the above formulas, it appears that both $\zeta_{\pm}$ define a conformal map from the two disks $\mathbb{D}_{\pm}$ (respectively) to a neighbourhood $\mathscr{D}_{n}$ of the origin which is homothetically expanding with $n^{\frac{3}{2}}$. The determination of the fractional root can be chosen so that $\zeta_{\pm}$ map the main arc $\Gamma_{m}$ (where $\operatorname{Re}\varphi=0$} and $\operatorname{Im}\varphi(z_{+})$ is decreasing from 0 at $a_{-}$ to $2{\rm i}\pi$ at $a_{+}$) to the negative real axis in the respective $\zeta_{\pm}$ planes, with the points $a_{\pm}$ being mapped to $\zeta_{\pm}=0$. The arc $\Gamma_{c}$ can be chosen so that it is mapped to the positive $\zeta_{\pm}$-axis while the arcs of $\mathscr{L}_{\pm}\cap\mathbb{D}_{\pm}$ are mapped to two straight segments with slopes $\pm 3\pi/2$.
Then, the jump matrices $J_{T_{0}}$ in the Riemann–Hilbert Problem 5.4 restricted to $\mathbb{D}_{+}$ can be rewritten as (we focus on the case of $\mathbb{D}_{+}$ for definiteness)
$J_{T_{0}}(z)=\begin{cases}\begin{bmatrix}1&Q(z){\rm e}^{-\frac{4}{3}\zeta_{+}^{\frac{3}{2}}}\\ 0&1\end{bmatrix},&z\in\Gamma_{c}\cap\mathbb{D}_{+},\vskip 2.84526pt\\ \begin{bmatrix}1&0\\ \displaystyle\frac{\displaystyle{\rm e}^{\frac{4}{3}{\zeta_{+}}^{\frac{3}{2}}}}{Q(z)}&1\end{bmatrix},&z\in\mathscr{L}_{\pm}\cap\mathbb{D}_{+},\vskip 2.84526pt\\ \begin{bmatrix}0&Q(z)\\ \displaystyle-\frac{1}{Q(z)}&0\end{bmatrix},&z\in\Gamma_{m}\cap\mathbb{D}_{+}.\end{cases}$ (5.7)
@@
$\mathcal{P}_{+}(z):={\bf A}\bigl(\zeta_{+}(z)\bigr)Q(z)^{-\frac{\sigma_{3}}{2}}{\rm e}^{\frac{2}{3}\zeta_{+}(z)^{\frac{3}{2}}\sigma_{3}}$
exhibits discontinuities across $(\Gamma_{c}\cup\mathscr{L}_{+}\cup\mathscr{L}_{-}\cup\Gamma_{m})\cap\mathbb{D}_{+}$ with jump matrices given exactly by (5.7), as a consequence of the RHP satisfied by the matrix ${\bf A}(\zeta)$ shown on Figure 11. Completely analogous expressions hold for $\mathcal{P}_{-}(z)$.